OpenStudy (anonymous):

Can anyone please help me with this:

OpenStudy (anonymous):

A tunnel is in the shape of a parabola. The maximum height is 16 m and it is 16 m wide at the base as shown below. A parabola opening down with vertex at the origin is graphed on the coordinate plane. The height of the parabola from top to bottom is 16 meters and its width from left to right is 16 meters. What is the vertical clearance 7 m from the edge of the tunnel?

OpenStudy (anonymous):

Please help me :)

OpenStudy (anonymous):

@ChristopherToni

OpenStudy (tkhunny):



OpenStudy (anonymous):

would it be 7?

OpenStudy (tkhunny):

What does " 7 m from the edge " mean?

OpenStudy (anonymous):

that it has to move from 8,0 on 7

OpenStudy (tkhunny):

You must find the equation of the parabola. \(y - 16 = Ax^{2}\) <== How to find the value of "A"?

OpenStudy (anonymous):

but how do i find a?

OpenStudy (anonymous):

would y be 16 because the height is 16?

OpenStudy (tkhunny):

I put the Origin in the middle of the road. This makes the Vertex (0,16). Put one of the other points (maybe (8,0)) in the equation and solve for A.

OpenStudy (anonymous):

im not understanding what im supposed to do

OpenStudy (tkhunny):

Do you believe \(y - 16 = Ax^{2}\)

OpenStudy (anonymous):

im not sure what you mean

OpenStudy (tkhunny):

We need the equation of that parabola. Is that it or not?

OpenStudy (anonymous):

well i use this : (x-h)^2=4p(y-k) that's for vertical and horizontal is : (y-k)^2=4p(x-h ) @tkhunny

OpenStudy (tkhunny):

Okay, that's fine. We know the vertex is (0,16). Does that help us along our way?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

What is the result after applying this information?

OpenStudy (anonymous):

well it's vertical so i would use:(x)^2=4p(y-16)

OpenStudy (tkhunny):

Does that look familiar? 1/A = 4p

OpenStudy (anonymous):

ohh yes the asymptote right?

OpenStudy (tkhunny):

No asymptote on a parabola. We don't have a Focus. We don't have a Directrix. How shall we find the value of p?

OpenStudy (anonymous):

by moving the (y-16) to the other side?

OpenStudy (tkhunny):

No good. We need another value. Just substitute (8,0) into that equation with "p" and see what happens.

OpenStudy (anonymous):

that's the focal width

OpenStudy (anonymous):

i got p=1

OpenStudy (tkhunny):

Keep in mind that p < 0 because this is a downward opening parabola. That MAY BE the focal width. Do we KNOW the Focus is at the Origin?

OpenStudy (tkhunny):

8^2=4p(0-16) 64 = 4*p*(-16) = -64p and we have p = -1

OpenStudy (tkhunny):

Okay, so what is the final equation?

OpenStudy (anonymous):

(x)^2=-4(y-16)

OpenStudy (tkhunny):

And for our final act, what value of y corresponds to x = 1?

OpenStudy (anonymous):

64=-4(y-16)

OpenStudy (tkhunny):

?? How does 1^2 = 64?

OpenStudy (anonymous):

i thought it was the 64 from the top from 8^2

OpenStudy (tkhunny):

Why would you think that? The equation is x^2=-4(y-16) Substitute x = 1.

OpenStudy (anonymous):

because idk where that 1 came from 1=-4(y-16)

OpenStudy (tkhunny):

Solve for y.

OpenStudy (anonymous):

can i distribute the -4?

OpenStudy (tkhunny):

If you want. Just find the value of y - any way you like.

OpenStudy (anonymous):

i got y=16

OpenStudy (tkhunny):

Note: It had better be a little less than 16.

OpenStudy (tkhunny):

1=-4(y-16) -1/4 = y-16 y = 16 - 1/4

OpenStudy (anonymous):

i moved it all to the other side... i finally got 15.75

OpenStudy (tkhunny):

There it is. Can you now answer the question?

OpenStudy (anonymous):

yes thank you for all of your help :)

OpenStudy (tkhunny):

Way to hang in there!! Understanding what the problem wanted was the first hard part. :-)

OpenStudy (anonymous):

yes but you really clarified that :D